OpenStudy (anonymous):

what multiplies to be -6 but adds to be 23

OpenStudy (anonymous):

-3 times -2 = -6 and 12 plus 11 =23 educated guess

OpenStudy (danjs):

23.258 and -0.258

OpenStudy (danjs):

aproximate

OpenStudy (danjs):

x + y = 23 and x*y = -6

OpenStudy (danjs):

\[\frac{ \sqrt{553}+23 }{ 2 } and \frac{ -\sqrt{553}-23 }{ 2 }\]

OpenStudy (danjs):

approx 23.258 and negative 0.258